AB INITIO TEST STUDY ON THE VERTICAL EXCITATION ENERGY OF COAr

\underline{MARY C. SALAZAR}, Department of Chemistry, Simon Bolivar University, Caracas 1080A, Venezuela; CARLOS MANZANARES I., Department of Chemistry and Biochemistry, Baylor University, Waco, TX 76798; HENRY J. CASTEJON, Department of Chemistry, Wilkes University, Wilkes-Barre PA 18766; and ANTONIO J. HERNANDEZ, Department of Chemistry, Simon Bolivar University, Caracas 1080A, Venezuela.

Accurate ab initio potential energy surfaces in the framework of the supermolecule approach are obtained using high-level ab initio methods of quantum chemistry. They are developed using perturbation methods and single and double excitation coupled-cluster theory with noniterative treatment of triple excitations CCSD(T) and the correlation-consistent basis sets supplemented by bond functions placed at the midpoint of the van der Waals bond. Vertical excitation electronic spectra of the diatomic CO − Ar van der Waals molecule will be presented.